OpenStudy (anonymous):

square root of 64 over square root of 3

OpenStudy (anonymous):



Parth (parthkohli):

\( \color{Black}{\Rightarrow\Large{\sqrt{64} \over \sqrt3} }\) \( \color{Black}{\Rightarrow \Large {8 \over \sqrt3}}\) \( \color{Black}{\Rightarrow \Large {8\sqrt3 \over3} }\)

OpenStudy (anonymous):

\[\frac{\sqrt{a}}{\sqrt{b}} = \sqrt{\frac{a}{b}}\]